OpenStudy (anonymous):

Please help, i need to find the derivative of y=1/(1+sec(x))^2.

OpenStudy (anonymous):

\[\frac{ d }{ dx }\frac{ 1 }{ (1+\sec(x))^2 }\] this?

OpenStudy (anonymous):

yes, essentially

OpenStudy (anonymous):

You can bring it up and use chain rule

OpenStudy (anonymous):

power + chain

OpenStudy (anonymous):

Give it an attempt and see what you get, then we'll see if you did it right or not.

OpenStudy (anonymous):

i know the chain rule, how do i bring it up?

OpenStudy (anonymous):

Exponent rules, if you remember \[\frac{ 1 }{ x^n } \implies x^{-n}\]

OpenStudy (anonymous):

omg... haha yes, ok, ill give it a show

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Yeah I think it's easiest if you bring it up in the numerator

OpenStudy (anonymous):

is it y'=tan(x)/4?

OpenStudy (anonymous):

Exponent rules, if you remember \[\left[ (1+\sec(x))^{-2} \right]' \implies -2(1+\sec(x))^{-3} \times (1+\sec(x))'\] did you get this far?

OpenStudy (anonymous):

i made an error in the exponent rule, one second

OpenStudy (anonymous):

Np, take your time

OpenStudy (anonymous):

Oh Idk why, but the lag removed my information above, w/e so you also have to keep in mind

OpenStudy (anonymous):

\[\frac{ d }{ dx } \sec(x) = \sec(x)\tan(x)\]

OpenStudy (anonymous):

my notes say \[d/dx(\sec(g(x))) = \sec(g(x))\tan(g(x))*g'(x)\] now i have \[y'=\frac{ \sec(x)\tan(x) }{ -2(1+\sec(x))^3 }\]

OpenStudy (anonymous):

is there anything more to simplify? if im at the right spot

OpenStudy (anonymous):

-2 should be in the numerator but other than that, you're good :)

OpenStudy (anonymous):

ahh ok, because its not to the -3rd power

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

Derivatives aren't so bad, it's just bunch of rules you remember and then just algebra

OpenStudy (anonymous):

thank you so much, honestly. and i know that, its exactly the fact that i cant remember a few of the rules, and just get blocked

OpenStudy (anonymous):

No worries :)

OpenStudy (anonymous):

A lot of people forget exponent rules especially so I'm going to put some up right now :P

OpenStudy (anonymous):

Glad my mistakes could help a few more :)